OpenStudy (anonymous):

Between which two integers does square root of 12 lie? A. 3 and 4 B. 5 and 7 C. 11 and 13 D. 143 and 145 @BloomLocke367

OpenStudy (bloomlocke367):

well, the best thing to do is make a number line.

OpenStudy (anonymous):

it is A

OpenStudy (anonymous):

well i learned this way from phi find the perfict square root of 12

OpenStudy (bloomlocke367):



OpenStudy (bloomlocke367):

etc..

OpenStudy (anonymous):

the answer is A

OpenStudy (bloomlocke367):

please don't give direct answers, it's against the CoC

OpenStudy (anonymous):

what is the coc

OpenStudy (anonymous):

1*1=1 2*2=4 3*3=9 4*4=16 so i guess it would be a because it is between 9 and 16 oh and it is code of conduct

OpenStudy (anonymous):

@unkownbeast911

OpenStudy (bloomlocke367):

yes, @blackdragon316 , it's between 3 and 4

OpenStudy (anonymous):

thanks but this wasnt one i needd help with i need one that is like 10times harder

OpenStudy (anonymous):

WHAT IS IT

OpenStudy (bloomlocke367):

well, open a new question and I'll try :)

OpenStudy (anonymous):

okay let me get it